# Big bounce and black bounce in quasi-topological gravity

### v1

@@ I Introduction
In the context of black holes, a parallel mechanism is proposed to resolve the singularity within the horizon of black holes. It suggests that infalling matter does not get crushed into a singularity but instead undergoes a quantum bounce at the Planck scale, and eventually re-emerges from a white hole. This process, often referred to as a black-to-white hole transition or a “black bounce”, leads to the formation of regular black holes. Numerous bouncing models have been investigated in the literature [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26].
@@ III Bounce in quasi-topological gravities
When $r<r_{0}$, the characteristic function diverges, and one fails to derive the equation of motion Eq. (7). This implies that the region $r<r_{0}$ is not part of the spacetime, and the metric is therefore regular everywhere. An interesting result is that the maximal value of the Kretschmann scalar $K^{2}\left|{}_{max}\right.\sim M^{2}/r_{0}^{2(D-1)}\sim\rho_{\text{Pl}}$, where $\rho_{\text{Pl}}$ is the Planck energy density, is independent of the mass $M$. Furthermore, the existence of a minimum value for $r$ distinguishes this metric from other regular metrics, such as the Hayward metric. It is, in fact, a bouncing solution: a particle in radial motion bounces upon reaching $r=r_{0}$ , instead of passing through the origin $r=0$.

### v2

@@ I Introduction
In the context of black holes, a parallel mechanism is proposed to resolve the singularity within the horizon of black holes. It suggests that infalling matter does not get crushed into a singularity but instead undergoes a quantum bounce at the Planck scale, and eventually re-emerges from a white hole. This process, often referred to as a black-to-white hole transition or a “black bounce”, leads to the formation of regular black holes. Numerous bouncing models have been investigated in the literature [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30].
@@ II Quasi-topological Gravities and Singularity Resolution
On the other hand, from the conservation equation for the stress-energy tensor of the fluid, one obtains
$\dot{\rho}+(D-1)(\rho+p)\frac{\dot{a}(\tau)}{a(\tau)}=0.$ (14)
Upon specifying an equation of state for the matter, the scale factor $a(t)$ can be obtained by solving above two equations.
@@ III Bounce in quasi-topological gravities
When $r<r_{0}$, the characteristic function diverges, and one fails to derive the equation of motion Eq. (7). This implies that the region $r<r_{0}$ is not part of the spacetime, and the metric is therefore regular everywhere. This result is expected, as noted in Refs. [19, 20], since in spherical symmetry there are no gravitational waves, and therefore a central potential must be generated by some distribution of matter. Furthermore, given that the energy density of matter field is bounded by the Planck scale $\rho\lesssim\rho_{\text{Pl}}$, where $\rho_{\text{Pl}}$ is the Planck energy density, a source of mass $M$ requires a matter distribution extending out to $r\sim\quantity(M/\rho_{\text{Pl}})^{1/3}\sim r_{0}$, forming a “Planck star”[9]. Consequently, a complete description of the spacetime solution must incorporate matter fields, for which the qOS model serves as a concrete example.
An interesting property of the metric (22) is that the maximal value of the Kretschmann scalar $K^{2}\left|{}_{max}\right.\sim M^{2}/r_{0}^{2(D-1)}\sim\rho_{\text{Pl}}$, is independent of the mass $M$. Furthermore, the existence of a minimum value for $r$ distinguishes this metric from other regular metrics, such as the Hayward metric. It is, in fact, a bouncing solution: a particle in radial motion bounces upon reaching $r=r_{0}$ , instead of passing through the origin $r=0$.